# Client Selection for Federated Policy Optimization with Environment Heterogeneity

### v5

@@ Abstract
The development of Policy Iteration (PI) has inspired many recent algorithms for Reinforcement Learning (RL), including several policy gradient methods that gained both theoretical soundness and empirical success on a variety of tasks. The theory of PI is rich in the context of centralized learning, but its study under the federated setting is still in the infant stage. This paper investigates the federated version of Approximate PI (API) and derives its error bound, taking into account the approximation error introduced by environment heterogeneity. We theoretically prove that a proper client selection scheme can reduce this error bound. Based on the theoretical result, we propose a client selection algorithm to alleviate the additional approximation error caused by environment heterogeneity. Experiment results show that the proposed algorithm outperforms other biased and unbiased client selection methods on the federated mountain car problem and the Mujoco Hopper problem by effectively selecting clients with a lower level of heterogeneity from the population distribution.
@@ 1 Introduction
Despite the significant progress of empirical works on FRL (Qi et al., 2021), the community’s understanding of FRL is still in its infancy, especially from the theoretical perspective. For example, the sample efficiency of Policy Gradient (PG) methods is typically low due to the large variance in gradient estimation. This issue could be exacerbated in the context of FL, where clients with heterogeneous environments can generate a diverse range of trajectories. To address this problem, a variance-reduced policy gradient method, namely Federated Policy Gradient with Byzantine Resilience (FedPG-BR), was proposed together with an analysis of the sample efficiency and convergence guarantee (Fan et al., 2021). While clients are assumed to be homogeneous in FedPG-BR, another line of work, termed FedKL (Xie and Song, 2023), noticed that the environment heterogeneity imposes an extra layer of difficulty in learning and proved that a Kullback-Leibler (KL) penalized local objective can generate a monotonically improving sequence of policies to accelerate convergence. The authors of QAvg & PAvg (Jin et al., 2022) provided convergence proof for the federated Q-Learning and federated PG. QAvg offered important insights regarding how the Bellman operators can be generalized to the federated setting and proposed a useful tool, i.e., the imaginary environment (the average of all clients’ environments), for analyzing FRL. However, there has not been any convergence analysis regarding Policy Iteration (PI) in FRL in the literature. Given PI’s application and theoretical importance, it is desirable to fill this knowledge gap and derive efficient FRL algorithms accordingly.
@@
Contributions. In this paper, we derive the error bound of Federated Approximate Policy Iteration (FAPI) under heterogeneous environments, which is not yet available in the literature. The derived error bound takes the heterogeneity level of clients into consideration and explicitly reveals its impact. Based on the error bound, we propose a client selection algorithm to improve the convergence speed of federated policy optimization. The efficacy of the proposed algorithm is validated on the federated mountain car and Mujoco Hopper problems.
@@ 2.1 Federated Reinforcement Learning
We model the local learning problem of each client as a finite-state infinite-horizon discounted Markov Decision Process (MDP). Accordingly, the FRL system consists of $N$ finite MDPs $\{(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)|n\in\{1,...,N\}\}$, where $\mathcal{S}$ denotes a finite set of states, $\mu$ represents the initial state distribution, $\mathcal{A}$ is a finite set of actions, and $\gamma\in(0,1)$ is the discount factor. The transition function $P_{n}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ represents the probability that the n-th MDP transits from state $s$ to $s^{\prime}$ after taking action $a$ (Sutton and Barto, 2018). The reward function $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ gives the expected reward for taking action $a$ in state $s$, and we assume rewards are bounded and non-negative. As a result, the $n$-th MDP $\mathcal{M}_{n}$ can be represented by a 6-tuple $(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)$ sharing the same state space, action space, initial state distribution, and reward function with other clients, but with possibly different transition probabilities. An client’s behavior is controlled by a stochastic policy $\pi_{n}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which outputs the probability of taking an action $a$ in a given state $s$. Furthermore, we define the state-value function
@@
where the expectation is performed over actions sampled from policy $\pi$ and states sampled from the transition probability $P_{n}$. It gives the expected return when the client starts from state $s$ and follows policy $\pi$ thereafter in the $n$-th MDP. In each round, every client aims to train a local policy to maximize its expected discounted reward
@@
$\max_{\pi}\eta(\pi)$ $=\mathbb{E}_{s_{0}\thicksim\mu}\left[\bar{V}^{\pi}(s_{0})\right].$ (3)
The above formulation covers both heterogeneous and homogeneous cases. In particular, the different MDPs, i.e., different transition probabilities, represent the heterogeneous environments experienced by different clients. All MDPs will be identical for the homogeneous case (Fan et al., 2021).
@@ 3 Error Bound of FAPI
In this section, we establish the error bound of FAPI under the framework of API and the imaginary MDP. The analysis shows that the FAPI process can be considered as learning in the imaginary MDP which eases the analysis. Without loss of generality, we focus our discussion on the tabular case (Watkins and Dayan, 1992; Singh et al., 2000; Jin et al., 2022), e.g., $\pi^{t+1}(a|s)=\sum_{n=1}^{N}q_{n}\pi^{t+1}_{n}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$ with full client participation. When function approximation is used, the error term introduced by the aggregation step can be easily incorporated into the analysis.
To establish the error bound of FAPI, we first examine the approximation error of policy evaluation and policy improvement within the FRL framework. By doing so, we can then derive the distance between the optimal value function and the value function produced by FAPI.
@@ 3.1 Federated Policy Evaluation
The following lemmas describe the relation between the averaged value function $\bar{V}^{\pi}$of policy $\pi$ and the value function $V_{I}^{\pi}$ of policy $\pi$ in the imaginary MDP.
@@
Readers are referred to Jin et al. (2022) for detailed proofs and discussions of Lemmas 3 and 4. Briefly, $V_{I}^{\pi}$ serves as a lower bound of $\bar{V}^{\pi}$. Among all policies, the optimal policy $\pi_{I}^{\ast}$ in the imaginary MDP is of particular interest since its value function $V_{I}^{\pi_{I}^{\ast}}=V_{I}^{\ast}$ is the largest lower bound of the average value function $\bar{V}^{\pi}$. Since there are approximation errors (7) in each client, we can only obtain an approximation $\bar{V}^{t}=\sum_{n}^{N}q_{n}V_{n}^{t}$ of the averaged value function $\bar{V}^{\pi^{t}}$ with
$\left\|\bar{V}^{t}-\bar{V}^{\pi^{t}}\right\|$ $\leq\sum_{n=1}^{N}q_{n}\delta_{n}=\bar{\delta}.$
Therefore, by the triangle inequality, we have
$\left\|\bar{V}^{t}-V_{I}^{\pi^{t}}\right\|$ $\leq\frac{\gamma R_{\max}\kappa_{1}}{(1-\gamma)^{2}}+\bar{\delta}=\dot{\delta}.$ (8)
@@ 3.2 Federated Policy Improvement
8:     The central server aggregates $V^{t}_{n}$ to obtain $\bar{V}^{t}$.
9:     Synchronize the global policy $\pi^{t}$ and value function $\bar{V}^{t}$ to every client.
@@
11:        Local update of client policy: $\left\|T^{\pi^{t+1}_{n}}_{n}\bar{V}^{t}-T_{n}\bar{V}^{t}\right\|\leq\epsilon_{n}$.
@@
14:     The central server performs aggregation: $\pi^{t+1}(a|s)\leftarrow\sum_{n=1}^{N}q_{n}\pi_{n}^{t+1}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$.
@@
9:     The central server performs aggregation: $\pi^{t+1}(a|s)\leftarrow\sum_{n=1}^{N}q_{n}\pi_{n}^{t+1}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$.
@@
$\left\|T_{I}^{\pi^{t+1}}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\frac{2\gamma R_{\max}\kappa_{1}}{1-\gamma}+\bar{\epsilon}=\epsilon^{\prime},\ \textnormal{where}\ \bar{\epsilon}=\sum_{n=1}^{N}q_{n}\epsilon_{n}.$
See Appendix A for the detailed proof.
@@
$\left\|T_{I}^{\pi^{t+1}}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\frac{2\gamma^{2}R_{\max}\kappa_{2}}{(1-\gamma)^{2}}+\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}+4\gamma\bar{\delta}+\bar{\epsilon}=\dot{\epsilon}.$ (9)
@@
$\left\|T_{I}^{\pi^{t+1}}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\left(\gamma+\gamma^{2}\right)R_{\max}\kappa_{m,n}}{(1-\gamma)^{2}}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}$
$\quad+2\gamma\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+2\gamma\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}=\hat{\epsilon}.$ (10)
@@
$\left\|T_{I}^{\pi^{t+1}}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\gamma R_{\max}\kappa_{m,n}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}=\acute{\epsilon}.$ (11)
@@ 3.3 Bounding the distance between value functions
where $\bar{V}^{\max}_{s}=\max\left\{\bar{V}^{\pi^{\ast}}(s),\bar{V}^{\pi^{\ast}_{I}}(s)\right\},\forall s\in\mathcal{S}$, and $\tilde{\epsilon}$ may be one of $\dot{\epsilon}$, ϵsuperscriptitalic-$\epsilon^{\prime}$, $\hat{\epsilon}$ or $\acute{\epsilon}$. More specifically, the error bound for Algorithm 2 with partial client participation is
@@
$\quad+\tilde{\mathcal{O}}\left(\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}\right),$ (13)

### v6

@@ Abstract
The development of Policy Iteration (PI) has inspired many recent algorithms for Reinforcement Learning (RL), including several policy gradient methods that gained both theoretical soundness and empirical success on a variety of tasks. The theory of PI is rich in the context of centralized learning, but its study under the federated setting is still in the infant stage. This paper investigates the federated version of Approximate PI (API) and derives its error bound, taking into account the approximation error introduced by environment heterogeneity. We theoretically prove that a proper client selection scheme can reduce this error bound. Based on the theoretical result, we propose a client selection algorithm to alleviate the additional approximation error caused by environment heterogeneity. Experiment results show that the proposed algorithm outperforms other biased and unbiased client selection methods on the federated mountain car problem, the Mujoco Hopper problem, and the SUMO-based autonomous vehicle training problem by effectively selecting clients with a lower level of heterogeneity from the population distribution.
@@ 1 Introduction
Despite the significant progress of empirical works on FRL (Qi et al., 2021), the community’s understanding of FRL is still in its infancy, especially from the theoretical perspective. For example, the sample efficiency of Policy Gradient (PG) methods is typically low due to the large variance in gradient estimation. This issue could be exacerbated in the context of FL, where clients with heterogeneous environments can generate a diverse range of trajectories. To address this problem, a variance-reduced policy gradient method, namely Federated Policy Gradient with Byzantine Resilience (FedPG-BR), was proposed together with an analysis of the sample efficiency and convergence guarantee (Fan et al., 2021). While clients are assumed to be homogeneous in FedPG-BR, another line of work, termed FedKL (Xie and Song, 2023), noticed that the environment heterogeneity imposes an extra layer of difficulty in learning and proved that a Kullback-Leibler (KL) penalized local objective can generate a monotonically improving sequence of policies to accelerate convergence. The authors of QAvg & PAvg (Jin et al., 2022) provided a convergence proof for the federated Q-Learning and federated PG. QAvg offered important insights regarding how the Bellman operators can be generalized to the federated setting and proposed a useful tool, i.e., the imaginary environment (the average of all clients’ environments), for analyzing FRL. More recently, FedSARSA (Zhang et al., 2024) studied the integration of FRL and SARSA, where SARSA is an on-policy Temporal Difference (TD) algorithm. However, there has not been any convergence analysis regarding Policy Iteration (PI) in FRL in the literature. Given PI’s application and theoretical importance, it is desirable to fill this knowledge gap and derive efficient FRL algorithms accordingly.
@@
Contributions. In this paper, we derive the error bound of Federated Approximate Policy Iteration (FAPI) under heterogeneous environments, which is not yet available in the literature. The derived error bound takes the heterogeneity level of clients into consideration and explicitly reveals its impact. Based on the error bound, we propose a client selection algorithm to improve the convergence speed of federated policy optimization. The efficacy of the proposed algorithm is validated on the federated mountain car problem, the Mujoco Hopper problems, and the SUMO-based autonomous vehicle training problem.
@@ 2.1 Federated Reinforcement Learning
We model the local learning problem of each client as a finite-state infinite-horizon discounted Markov Decision Process (MDP). Accordingly, the FRL system consists of $N$ finite MDPs $\{(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)|n\in\{1,...,N\}\}$, where $\mathcal{S}$ denotes a finite set of states, $\mu$ represents the initial state distribution, $\mathcal{A}$ is a finite set of actions, and $\gamma\in(0,1)$ is the discount factor. The transition function $P_{n}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ represents the probability that the n-th MDP transits from state $s$ to $s^{\prime}$ after taking action $a$ (Sutton and Barto, 2018). The reward function $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ gives the expected reward for taking action $a$ in state $s$, and we assume rewards are bounded and non-negative. As a result, the $n$-th MDP $\mathcal{M}_{n}$ can be represented by a 6-tuple $(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)$ sharing the same state space, action space, initial state distribution, and reward function with other clients, but with possibly different transition probabilities. A client’s behavior is controlled by a stochastic policy $\pi_{n}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which outputs the probability of taking an action $a$ in a given state $s$. Throughout this work, we consider parameterized policies $\pi^{\theta}$ where $\pi^{\theta}(a|s)$ is a differentiable function of the parameter vector $\theta$. For parameterized policies with t-indexed notation, we omit the parameter vector for notation simplicity and write5 $\pi^{t}$ and $\pi^{t}_{n}$ for the global policy $\pi^{\theta^{t}}$ and the $n$-th local policy $\pi^{\theta^{t}_{n}}$, respectively. Furthermore, we define the state-value function
@@
where the expectation is performed over actions sampled from policy $\pi$ and states sampled from the transition probability $P_{n}$. It gives the expected return when the client starts from state $s$ and follows policy $\pi$ thereafter in the $n$-th MDP. For parameterized value functions with t-indexed notation, we omit the parameter vector $w$ for notation simplicity and write $V^{t}$ and $V^{t}_{n}$ for $V^{w^{t}}$ and $V^{w^{t}_{n}}$, respectively. In each round, every client aims to train a local policy to maximize its expected discounted reward
@@
$\max_{\pi}\eta(\pi)\quad\text{where}\quad\eta(\pi)$ $=\mathbb{E}_{s_{0}\thicksim\mu}\left[\bar{V}^{\pi}(s_{0})\right].$ (3)
The above formulation covers both heterogeneous and homogeneous cases. In particular, the different MDPs, i.e., different transition probabilities, represent the heterogeneous environments experienced by clients. All MDPs will be identical for the homogeneous case (Fan et al., 2021). It is worth noting that the optimization problem in (3) is often referred to as the Weighted Value Problem (WVP) in the latent MDP literature. Finding the optimal solution of WVP is NP-hard (Steimle et al., 2021). In contrast, an error bound showing the distance between the obtained policy and the optimal policy is feasible as demonstrated in Section 3.
@@ 3 Error Bound of FAPI
In this section, we establish the error bound of FAPI under the framework of API and the imaginary MDP. The analysis shows that the FAPI process can be considered as learning in the imaginary MDP which eases the analysis. We focus our discussion on policy with function approximation where the error term introduced by the aggregation step is nontrivial.
To establish FAPI’s error bound, we first examine the approximation error of policy evaluation and improvement within the FRL framework. Then, we can derive the distance between the optimal value function and the value function produced by FAPI.
@@ 3.1 Federated Policy Evaluation
The following lemmas describe the relation between the averaged value function $\bar{V}^{\pi}(s)=\sum_{n=1}^{N}q_{n}V_{n}^{\pi}(s),\forall s\in\mathcal{S}$ of policy $\pi$ and the value function $V_{I}^{\pi}$ of policy $\pi$ in the imaginary MDP.
@@
Readers are referred to Jin et al. (2022) for detailed proofs and discussions of Lemmas 3 and 4. Briefly, $V_{I}^{\pi}$ serves as a lower bound of $\bar{V}^{\pi}$. Among all policies, the optimal policy $\pi_{I}^{\ast}$ in the imaginary MDP is of particular interest since its value function $V_{I}^{\pi_{I}^{\ast}}=V_{I}^{\ast}$ is the largest lower bound of the average value function $\bar{V}^{\pi}$. Since there are approximation errors (7) in each client and aggregation error in the central server, we can only obtain an approximation $V^{t}$ of the averaged value function $\bar{V}^{\pi^{t}}$ with
$\left\|V^{t}-\bar{V}^{\pi^{t}}\right\|\leq\left\|V^{t}-\bar{V}^{t}\right\|+\left\|\bar{V}^{t}-\bar{V}^{\pi^{t}}\right\|$ $\leq\bar{\varepsilon}_{w}+\sum_{n=1}^{N}q_{n}\delta_{n}=\bar{\delta},$
where $\bar{V}^{t}(s)=\sum_{n}^{N}q_{n}V_{n}^{t}(s),\forall s\in\mathcal{S}$ is the average of the local approximation and $\bar{\varepsilon}_{w}=\max_{t}\left\|V^{t}-\bar{V}^{t}\right\|$ is the error induced by value function aggregation. Note that we postpone the discussion of the error $\varepsilon_{w}$ induced by value function aggregation to the end of this section. Therefore, by the triangle inequality, we have
$\left\|V^{t}-V_{I}^{\pi^{t}}\right\|$ $\leq\frac{\gamma R_{\max}\kappa_{1}}{(1-\gamma)^{2}}+\bar{\delta}=\dot{\delta}.$ (8)
@@ 3.2 Federated Policy Improvement
8:     The central server aggregates $V^{t}_{n}$ to obtain $V^{t}$: $w^{t+1}\leftarrow\sum_{n=1}^{N}q_{n}w_{n}^{t+1}$.
9:     Synchronize the global policy $\pi^{t}$ and value function $V^{t}$ to every client.
@@
11:        Local update of client policy: $\left\|T^{\pi^{t+1}_{n}}_{n}V^{t}-T_{n}V^{t}\right\|\leq\epsilon_{n}$.
@@
14:     The central server aggregates $\pi^{t+1}_{n}$ to obtain $\pi^{t+1}$: $\theta^{t+1}\leftarrow\sum_{n=1}^{N}q_{n}\theta_{n}^{t+1}$.
@@
9:     The central server aggregates $\pi^{t+1}_{n}$ to obtain $\pi^{t+1}$: $\theta^{t+1}\leftarrow\sum_{n=1}^{N}q_{n}\theta_{n}^{t+1}$.
@@
$\left\|T_{I}^{\pi^{t+1}}V^{t}-T_{I}V^{t}\right\|\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+\frac{2\gamma R_{\max}\kappa_{1}}{1-\gamma}+\bar{\epsilon}=\epsilon^{\prime},$
where $\bar{\epsilon}=\sum_{n=1}^{N}q_{n}\epsilon_{n}$ and $\bar{\varepsilon}_{\theta}=\max_{t>0,s\in\mathcal{S}}\left\|\pi^{t+1}(\cdot|s)-\sum_{n=1}^{N}q_{n}\pi^{t+1}_{n}(\cdot|s)\right\|_{2}$.
See Appendix A for the detailed proof. Note that we postpone the discussion of the error $\varepsilon_{\theta}$ induced by policy aggregation to the end of this section.
@@
$\left\|T_{I}^{\pi^{t+1}}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\frac{2\gamma^{2}R_{\max}\kappa_{2}}{(1-\gamma)^{2}}+\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}+4\gamma\bar{\delta}+\bar{\epsilon}=\dot{\epsilon}.$ (9)
@@
$\left\|T_{I}^{\pi^{t+1}}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\left(\gamma+\gamma^{2}\right)R_{\max}\kappa_{m,n}}{(1-\gamma)^{2}}$
$\quad+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+2\gamma\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+2\gamma\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}=\hat{\epsilon},$ (10)
where $\tilde{\varepsilon}_{\theta}=\max_{t>0,s\in\mathcal{S}}\left\|\pi^{t}(\cdot|s)-\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t}_{m}(\cdot|s)\right\|_{2}$.
@@
$\left\|T_{I}^{\pi^{t+1}}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\gamma R_{\max}\kappa_{m,n}}{1-\gamma}$
$\quad+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}=\acute{\epsilon}.$ (11)
@@ 3.3 Bounding the distance between value functions
where $\bar{V}^{\max}_{s}=\max\left\{\bar{V}^{\pi^{\ast}}(s),\bar{V}^{\pi^{\ast}_{I}}(s)\right\},\forall s\in\mathcal{S}$, and $\tilde{\epsilon}$ may be one of $\dot{\epsilon}$, $\epsilon^{\prime}$, $\hat{\epsilon}$ or $\acute{\epsilon}$. More specifically, the error bound for Algorithm 2 with partial client participation is
@@
$\quad+\tilde{\mathcal{O}}\left(\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}+\bar{\varepsilon}_{w}+\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}\right),$ (13)
@@
3.4 Impact of Aggregation Error
We have defined three terms to quantify the impact of policy aggregation and value function aggregation, i.e., $\tilde{\varepsilon}_{\theta}$, $\bar{\varepsilon}_{\theta}$ and $\bar{\varepsilon}_{w}$. To further investigate how these errors affect the convergence of FAPI, there are two possible approaches: 1) Performing a general analysis with a few standard assumptions in the convex optimization literature; and 2) Carrying out the analysis with a specific neural network parameterization to gain insight into network configuration that can affect the approximation error. In light of the recent breakthrough in overparameterized neural networks (Arora et al., 2019; Cai et al., 2019; Wang et al., 2019; Liu et al., 2019), we employ the second strategy. To that end, we analyze the impact of aggregation error with the following two-layer ReLU-activated neural networks to parameterize the state value function and policy, respectively, as
$u_{w}(s)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\sigma(w_{i}^{T}(s)),$ (14)
$f_{\theta}(s,a)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\sigma(\theta_{i}^{T}(s,a)),$ (15)
where $m$ is the width of the network, $\theta=(\theta_{i},\dots\theta_{m})^{T}\in\mathcal{R}^{m(d_{s}+d_{a})}$, $w=(w_{i},\dots w_{m})^{T}\in\mathcal{R}^{md_{s}}$ denotes the input weights, and $b_{i}\in\{-1,1\},\forall i\in[m]$ is the output weight. Without loss of generality, we assume that $a\in\mathbb{R}^{d_{a}},s\in\mathbb{R}^{d_{s}}$ and denote $d=d_{s}+d_{a}$. Given arbitrary $\hat{R}_{\vartheta}<\infty$, we consider the following parameter initialization, where $\vartheta$ is a placeholder for $\theta$ and $w$:
$\mathbb{E}_{\text{init}}\left[\vartheta^{0}_{i,j}\right]=0,\mathbb{E}_{\text{init}}\left[\left(\vartheta^{0}_{i,j}\right)^{2}\right]=\frac{1}{d\cdot m},\forall i\in[m],j\in[d],$
$b_{i}\thicksim\text{Unif}(\{-1,1\}),\mathbb{E}_{\text{init}}\left[\left\|\vartheta^{0}_{i}\right\|_{2}^{-2}\right]<\infty,\left\|\vartheta^{0}_{i}\right\|_{2}\leq\hat{R}_{\vartheta},\forall i\in[m].$ (16)
In other words, $0<\left\|\vartheta^{0}_{i}\right\|_{2}\leq\hat{R}_{\vartheta},\forall i\in[m]$. The output weights $b_{i},\forall i\in[m]$ are fixed. The input weights $\vartheta_{i},\forall i\in[m]$ are projected into a ball centered at the initial parameter, i.e., $\theta\in\mathcal{B}^{0}_{\mathcal{R}_{\theta}}=\left\{\theta:\left\|\theta-\theta^{0}\right\|_{2}\leq\mathcal{R}_{\theta}\right\},w\in\mathcal{B}^{0}_{\mathcal{R}_{w}}=\left\{w:\left\|w-w^{0}\right\|_{2}\leq\mathcal{R}_{w}\right\}$. Then, the state value function and Softmax policy are parameterized by
$V^{t}(s)=u_{w^{t}}(s),\pi^{t}(a|s)=\frac{\exp(f_{\theta^{t}}(s,a))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(f_{\theta^{t}}(s,a^{\prime}))},\forall a\in\mathcal{A},s\in\mathcal{S}.$ (17)
The following lemma quantifies the aggregation error under the above setting.
Lemma 12
By utilizing the initialization scheme in (16), the policy parameterization in (17), and the neural network parameterization in (14) and (15), we have
$\bar{\varepsilon}_{w}$ $=\mathcal{O}\left(R_{w}^{6/5}m^{-1/10}\hat{R}_{w}^{2/5}\right),$ (18)
$\tilde{\varepsilon}_{\theta}$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right),$ (19)
$\bar{\varepsilon}_{\theta}$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right).$ (20)
See Appendix G for the detailed proof.
Remark 13
Lemma 12 indicates that the aggregation error is determined by the neural network parameterization ($m$) and optimization method ($R_{\theta}$ and $R_{w}$). Moreover, it can be observed from the proof that the error of value function aggregation is minimized when using linear function approximation, e.g., linear regression and tabular implementation, and the error of policy aggregation is minimized when using log-linear policy parameterization, e.g., Softmax policy with linear regression. In other words, there is a trade-off between training (non-linearity) and aggregation error (linearity). A non-linear function approximation can fit the training data well, but suffers from high aggregation error. In contrast, the linear function approximation may not be able to solve complex problems, but it does not introduce aggregation error.